OpenStudy (anonymous):

Help Please! (Answer will receive metal) Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. A. Passes through (-1, 4) and (-5, 2) B. Slope 2, y-intercept -4 C. Has an x intercept of 6 and a y-intercept of 3. D. passes through (1, 2) with slope -5/3 Thanks so much <3

OpenStudy (anonymous):

assuming point 1 as (x1,y1) and point 2 as (x2,y2), slope=(y2-y1)/(x2-x1), solve for c in y=mx+c by substituting both points, here m is slope and c is y intercept for option c, points are (6,0) and (0,3) solve for m and c If u have any doubts regarding substitutions, ping me

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

can you help? @CO_oLBoY

OpenStudy (anonymous):

@Mclovin21

OpenStudy (anonymous):

Thank you (: